# Increasing Trust in Language Models through the Reuse of Verified Circuits

### v7

@@ Abstract
Language Models (LMs) are increasingly used for a wide range of prediction tasks, but their training can often neglect rare edge cases, reducing their reliability. Here, we define a stringent standard of trustworthiness whereby the task algorithm and circuit implementation must be verified, accounting for edge cases, with no known failure modes. We show that a model can be trained to meet this standard if built using mathematically and logically specified frameworks. In this paper, we fully verify an auto-regressive transformer model that performs n-digit integer addition. To exhibit the reusability of verified modules, we insert the trained integer addition model into a larger untrained model and train the combined model to perform both addition and subtraction. We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model. We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of language models built using them. The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety and interpretability of language models.
@@
Philip Quirke Apart Research. philipquirkenz@gmail.com                        Clement Neo Apart Research.                        Fazl Barez Apart Research. Department of Engineering Sciences, University of Oxford fazl@robots.ox.ac.uk
@@ 1 Introduction
In creating and training a model, we would ideally want the model to have a very high standard of accuracy and trustworthiness. We aim to achieve this by holding the model to a standard that we term known-good. We define a model performing a task to be known-good if:
@@
Exhaustive testing of a model task may be infeasible. For instance, when adding two 5-digit integers (e.g. 12345+67890) there are ten billion variations. Some tasks can be conceptualized within an existing formal framework that allows identification of all edge cases the model must handle. For example, in 5-digit addition, the most uncommon edge case is 55555+44445=100000, which requires a carry bit to cascade through all digits, occurring in only 0.002% of cases. A known-good model must incorporate algorithms to manage all known edge cases. A known-good model would hence have verified [^1] In this paper, ’verified’ has the mechanistic interpretability meaning that a specific group of interconnected neurons within a neural network reliably and causally contributes to the model’s output in a meaningful, understandable way, with supporting empirical evidence. circuits that perform the task accurately.
In this paper, we detail the development and interpretation of a known-good model for addition. Our findings indicate that the model constructs a specific circuit for each edge case, with these circuits sharing intermediate results. We confirm the validity of the entire set of circuits, ensuring they cover all identified edge cases. The model achieves a very low training loss and has six nines (99.9999%) accuracy. The model hence achieves our standard of ’known-good’. Additionally, we develop a “mixed" model capable of both addition and subtraction, incorporating the known-good addition model. This mixed model has six nines accuracy, and extensively reuses the addition circuits for both operations and learns new circuits, facilitating the interpretation of the model’s algorithm. This approach marks progress toward a known-good model for both addition and subtraction.
@@
  • Demonstrating a proof of concept for re-using a known good model in the training of another larger, more-capable model, simplifying the interpretation of the more-capable model’s algorithm.
@@ 3 Methodology
A transformer model may learn an addition algorithm that differs significantly from the traditional human process. Here we define an alternative mathematical framework for addition, showing it is equivalent to the traditional process and so is mathematically correct. Then we show that our model implements this alternative framework.
@@ 3.1 Mathematical Framework
First, the model will need to sum pairs of digits. As in Quirke and Barez (2024), we define the “Simple Addition" sub-task $A_{n}.SA$, which naively calculates the sum of pairs of question digits, as:
@@
When there is no carry bit from the previous digit $A_{n}=A_{n}.SA$. The model will also need learn how to calculate whether the addition creates a carry bit. As in Quirke and Barez (2024), we define the "Simple Carry" sub-task $A_{n}.SC$:
@@
The model in Quirke and Barez (2024) is limited to the framework described so far. While it is capable of handling simple addition and carry bits generated directly from the addition of digit pairs, it encounters difficulties when faced with problems involving ’cascading carry’ bits, wherein a carry bit from one digit position propagates to the next.
To illustrate this limitation, consider the problem "00144+00056=000210". When adding the digits in the tens position, 4+5, no carry bit is generated directly. However, due to the carry bit propagating from the ones position, a carry bit is ultimately required in the tens position. A model that solely calculates the sums of digit pairs and considers carry bits resulting from the direct addition of those pairs will fail to accurately resolve this type of problem. In such cases, it would yield an incorrect result, such as "144+56=110". The 1-layer model could cascade the carry bit up to two digits across, but not more.
@@ 3.2 Extending the Mathematical Framework
To answer "44444+55556=" with "100000", an accurate model must predict the first answer digit A5 as "1". To do so, an accurate model must implement a “carry one cascade" circuit, which combines the “carry one" information from all five digits. This is especially hard as the model predicts answer tokens from left to right.
@@
Figure 3 diagrams how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition. This algorithm is like the 99% accurate Quirke and Barez (2024) algorithm but contains an additional circuit (the shaded boxes) to calculate “cascading carry one" data to predict with 99.9999% accuracy.
@@ 3.3 Techniques
    Question Complexity. We categorized questions by the complexity of the computation required to solve the question (Refer App. G). The addition categories (named S0 to S4) reflect the number of sequential digits that a “carry one" cascades through. The subtraction categories (named M0 to M4) reflect the number of sequential digits that a “borrow one" cascades through. We analysed which nodes were needed for correct prediction of each category.
@@ 4.1 Training a Five-Digit Addition Model
The Quirke and Barez (2024) 5-digit 1-layer addition model achieved an accuracy of $\sim$ 99%. Our experiments suggested that a 2-layer, 3-head model was the smallest configuration capable of achieving 99.9999% accuracy (see App. D for alternatives tested). This configuration effectively doubled the computational power compared to the 1-layer model (see App. B for mode model configuration details). Moreover, a 2-layer model introduces the capability to “compose" the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
@@
4.2 Investigating Five-Digit Addition
Ablation experiments targeting the nodes revealed that the model depends only on nodes located in nine token positions (Figs 6 and 7). Further ablation experiments show that for these nine token positions, model uses 36 nodes in predictions. The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 5 and Table  1), providing insight into the specific computations performed at each node. For each answer digit $A_{n}$, using test questions corresponding to the ST8, ST9 and ST10 categories, we performed PCA on the nodes yielding interpretable results. Specifically nine “node and answer-digit” combinations (see Figs 8 and  9) showed strong clustering of the questions aligned to the ST8, ST9 and ST10 categories.
@@ 4.2 Investigating Five-Digit Addition
Table 1: The nodes used in predictions across token positions. For all addition and mixed models studied, before the "=" token, each node often calculates data used to predict multiple answer digits. After the "=" token, all nodes in a given token position are used to predict a single answer digit. For this sample model, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5.
@@
For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes. Using ablation intervention to override each node, we verified each node’s role in executing the expected algorithmic step. Figure 3 diagrams the addition algorithm, while App. J details each nodes purpose for this model instance. The model adheres to all known constraints. Empirically, it achieves six nines accuracy. In summary, we concluded that this specific instance of the 5-digit addition model documented above has six nines accuracy, is well-understood and well-functioning and hence known-good.
@@
4.3 Training n-digit Addition models
4.4 Investigating n-digit Addition models
We implemented a declarative way to describe each sub-task. For example, a sub-task can be described as primarily attending to question digits D2 and D’2, influencing answer digit A3, impacting S0 but not S1 complexity questions, with a specific PCA result, and a specific ablation test. Using the declaration, we searched for and identified nodes performing all sub-tasks in all the models.
An algorithm hypothesis (e.g the one described in section 4.2) is then described in terms of the sub-tasks that must exist in the model and their inter-relationships. For example, our addition algorithm description states (in part) that the model must implement $A_{n}.ST$ sub-tasks for each question digit before the “=” token. We implemented an n-digit algorithm hypothesis declaration framework and a way to test the declaration against the sub-tasks found in each model and map the results. (Table 2 shows an example map.)
@@ 4.4 Investigating n-digit Addition models
We found that our seven 2-layer addition models all implement our addition algorithm. Given these seven models have six nines accuracy, and all implement the same valid algorithm, we can state these models are known-good.
@@ 4.6 Investigating n-digit Mixed models
Table 3: Mixed models re-use most inserted addition-model nodes. Many inserted nodes become polysemantic during training. That is they perform addition (S), positive-answer subtraction (M) and negative-answer subtraction (N) sub-tasks simultaneously. For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
@@ 4.8 Mixed model summary
The 7 mixed experimental models all have five or six nines accuracy. They all contain the same sub-tasks - suggesting they all implement the same algorithm. As we don’t fully understand the algorithm yet, we can’t assert that these mixed models are known-good.
@@ 5 Conclusion
In this paper, we successfully trained 5-, 6- and 10-digit, 2-layer, 3-head addition models that we have verified as known-good. They all implement the same algorithm, albeit with some implementation variations.
A significant aspect of our work was demonstrating component reuse. We achieved this by integrating an existing addition model into a new larger “mixed” model designed for both addition and subtraction tasks. This mixed model has six nines accuracy.
This integration allowed us to leverage our in-depth understanding of the addition model’s algorithm to unravel the workings of the mixed model’s algorithm. Interestingly, the mixed model re-uses most of the inserted addition model nodes, and “upgrades” most of them to become polysemantic - performing the inserted addition subtask and the corresponding subtraction subtasks simultaneously.
This paper lends support to the assertion by (Michaud et al., 2023) that many natural prediction problems can be broken down into a finite set of “quanta” computations. These quanta are essential for models to learn to minimize loss. Furthermore, it aligns with the concept that comprehending a network’s functionality is fundamentally about identifying and understanding its sub-quanta.
@@ 6 Limitations and Future Work
Our research opens up several promising avenues for future exploration. A key objective is to complete the reverse-engineering of the mixed model’s algorithm. We need to understand how the polysemantic nodes and the OPR+SGN node outputs are represented in the residual stream and show the MLP layers combine this information to select the appropriate question class data to accurately predict the answer digit.
Additionally, we aim to develop a comprehensive model that proficiently handles integer n-digit addition, subtraction, and multiplication, ensuring its status as known-good.
We believe this approach can be re-used to create known-good models in fields like logical reasoning (for AND, OR, Converse, Inverse etc tasks) and planning.
Using a smaller model to improve an existing Large Language Model (LLM) trustworthiness and capabilities is an active area of research. Approaches include composing multiple models (Bansal et al., 2024), LM up-scaling which essentially emulates fine-tuning a large pre-trained model using a small model (Mitchell et al., 2023), and this paper’s insertion of an accurate model into an untrained model. Other research (Voita et al., 2023; Hu et al., 2021) indicates the presence of “spare” neurons in LLMs and the potential for integrating small-scale modifications to fix erroneous circuits in LLMs. Further research into the benefits of using known-good models to increase LLM trustworthiness and capabilities is justified.
Moreover, developing methods to incorporate compact known-good models into LLMs, whether untrained or trained, could democratize AI Safety research. It enables small teams to focus on specific functional areas, crafting component models to a defined quality standard for integration into various LLMs.
Our automated framework, for discovering the location of algorithm sub-tasks in models, sped up our research. Our framework can be re-used to ensure that known-good components retain their functionality when integrated into LLMs.
Some of our automated framework is specific to our maths models - much is generic and reuseable on other types of models. Our automation focuses more on attention heads and less on MLP layers. Extending our framework to integrate the output data from complimentary frameworks (e.g. ACDC’s node connectivity strength data (Conmy et al., 2023)) would provide more automated insights into a model’s behavior, further speeding up research.
We made progress on a declarative language to describe an algorithm in terms of the number and types sub-tasks that must exist in a model for the algorithm to feasibly explain the model’s implementation. We made progress on a framework to automatically test an algorithm described in this way against a specific model. Further work is required to fully realize its potential.
@@ Appendix E Appendix: Experimental models
Num Digits Num Layers Num Heads Training Steps Seed Addition / Sub Accuracy Training loss Heads used MLPs used
@@
5 1 3 30K 372001 two nines 9.4e-2 15 6
5 2 3 15K 372001 six nines 1.6e-8 30 16
5 2 3 40K 372001 six nines 2.0e-9 22 15
6 2 3 15K 372001 six nines 1.7e-8 31 17
6 2 3 20K 173289 six nines 1.5e-8 28 17
6 2 3 20K 572091 six nines 7.0e-9 35 17
6 2 3 40K 372001 six nines 2.0e-9 29 17
10 2 3 20K 572091 six nines 7.0e-9 44 28
@@
6 2 3 40K 572091 six / five nines 2.4e-8 57 21
6 3 3 40K 572091 six / five nines 1.8e-8 70 35
6 3 3 80K 572091 six / five nines 1.6e-8 75 35
6 3 4 40K 372001 six / six nines 8.0e-9 72 26
6 3 4 40K 173289 five / five nines 1.4e-8 60 29
6 3 4 50K 572091 six / five nines 2.9e-8 79 29
10 3 3 50K 572091 five / five nines 6.3e-7 90 45
@@
6 4 3 40K 372001 four / two nines 3.0e-4 115 53
Table 6: Main experimental models studied

### v8

@@ Abstract
Language Models (LMs) are increasingly used for a wide range of prediction tasks, but their training can often neglect rare edge cases, reducing their reliability. Here, we define a stringent standard of trustworthiness whereby the task algorithm and circuit implementation must be verified, accounting for edge cases, with no known failure modes. We show that a model can be trained to meet this standard if built using mathematically and logically specified frameworks. In this paper, we fully verify an auto-regressive transformer model that performs n-digit integer addition. To exhibit the reusability of verified modules, we insert the trained integer addition model into a larger untrained model and train the combined model to perform both addition and subtraction. We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model. We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of LMs built using them. The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety and interpretability of LMs.
@@
Philip Quirkethanks:    Core research contributor., Clement Neo†‡, Fazl Barezfootnotemark: †‡ Apart Research, University of Oxford
@@ 1 Introduction
In creating and training a model, we aim for high accuracy and trustworthiness. We achieve this by holding the model to a standard we term known-good. We define a model performing a task to be known-good if:
@@
Exhaustive testing of a model task may be infeasible. For instance, when adding two 5-digit integers (e.g. 12345+67890) there are ten billion variations. Some tasks can be conceptualized within an existing formal framework that allows identification of all edge cases the model must handle. For example, in 5-digit addition, the most uncommon edge case is 55555+44445=100000, which requires a carry bit to cascade through all digits, occurring in only 0.002% of cases. A known-good model must incorporate algorithms to manage all known edge cases. A known-good model must have verified [^1] In this paper, ’verified’ has the mechanistic interpretability meaning that a specific group of interconnected neurons within a neural network reliably and causally contributes to the model’s output in a meaningful, understandable way, with supporting empirical evidence. circuits that perform the task accurately.
In this paper, we detail the development and interpretation of a known-good model for addition. Our findings indicate that the model constructs a specific circuit for each edge case, with these circuits sharing intermediate results. We confirm the validity of the entire set of circuits, ensuring they cover all identified edge cases. The model achieves a very low training loss and has six nines (99.9999%) accuracy. The model hence achieves our known-good standard. Additionally, we develop a “mixed" model capable of both addition and subtraction, incorporating the known-good addition model. This mixed model has six nines accuracy, and extensively reuses the addition circuits for both operations, facilitating the interpretation of the model’s algorithm. We make progress toward a known-good model for both addition and subtraction.
@@
  • Demonstrating a proof of concept for re-using a known good model in the training of another larger, more-capable model, simplifying the interpretation of the new model’s algorithm.
@@ 3 Methodology
Transformer models may learn addition algorithms different from traditional human methods. We define an alternative, mathematically-equivalent framework for addition and demonstrate our model implements this approach.
@@ 3.1 Mathematical Framework
First, we adopt the framework from Quirke and Barez (2024) for our model’s addition process. The “Simple Addition" sub-task $A_{n}.SA$, which naively calculates the sum of digit pairs, is defined as:
@@
When there is no carry bit from the previous digit $A_{n}=A_{n}.SA$. The “Simple Carry" sub-task $A_{n}.SC$ determines whether the addition creates a carry bit:
@@
While Quirke et al.’s model is capable of handling simple addition and carry bits generated directly from digit pair addition, it encounters difficulties with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
Consider “00144+00056=000210". Adding 4+5 in the tens position doesn’t generate a carry bit directly, but a carry bit propagates from the ones position. A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110". The Quirke et al. 1-layer model could cascade carry bits across two digits, but not three or more.
@@ 3.2 Extending the Mathematical Framework
To answer “44444+55556=" with “100000", an accurate model must predict the first answer digit A5 as “1". To do so, an accurate model must implement a “carry one cascade" circuit, which combines the “carry one" information from all five digits. This is especially hard as the model predicts answer tokens from left to right.
@@
Figure 3 diagrams how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition. This algorithm is like the 99% accurate Quirke et al. algorithm but contains an additional circuit (the shaded boxes) to calculate “cascading carry one" data to predict with 99.9999% accuracy.
@@ 3.3 Techniques
    Question Complexity. We categorized questions by computational complexity (App. G). Addition categories (S0-S4) and subtraction categories (M0-M4) reflect the number of sequential digits a “carry one" or “borrow one" cascades through, respectively. We analyzed which nodes were necessary for correct predictions in each category.
@@ 4.1 Training a Five-Digit Addition Model
The Quirke et al. 5-digit 1-layer addition model achieved an accuracy of $\sim$ 99%. Our experiments suggested that a 2-layer, 3-head model was the smallest configuration capable of achieving 99.9999% accuracy (see App. D for alternatives tested). This configuration effectively doubled the computational power compared to the 1-layer model (see App. B for mode model configuration details). Moreover, a 2-layer model introduces the capability to “compose" the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
@@
Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact. Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5. For all addition and mixed models studied, before the "=" token, each node often calculates data used to predict multiple answer digits. After the "=" token, all nodes in a given token position are used to predict a single answer digit.
4.2 Investigating Five-Digit Addition
Ablation experiments targeting the nodes revealed that the model depends only on nodes located in nine token positions (Figs 6 and 7). Further ablation experiments show that for these nine token positions, the model uses 36 nodes in predictions. The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 5 and Table  1), providing insight into the specific computations performed at each node. For each answer digit $A_{n}$, using test questions corresponding to the ST8, ST9 and ST10 categories, we performed PCA on the nodes yielding interpretable results. Specifically nine “node and answer-digit” combinations (see Figs 8 and  9) showed strong clustering of the questions aligned to the ST8, ST9 and ST10 categories.
@@ 4.2 Investigating Five-Digit Addition
For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes and verified each node’s role using ablation intervention. Figure 3 diagrams the algorithm, with node details in App. J. The model adheres to all known constraints and achieves six nines accuracy. We concluded this model instance is well-understood, well-functioning, and hence known-good.
@@
4.3 Training n-digit Addition models
4.4 Investigating n-digit Addition models
We developed a declarative method to outline each sub-task. For instance, a sub-task may focus on question digits D2 and D’2, affect answer digit A3, influence S0 but not S1 complexity questions, and have specific PCA results and ablation tests. Using this declaration, we identified nodes performing these sub-tasks across all models.
An algorithm hypothesis, such as the one in section 4.2, is described by the required sub-tasks and their relationships. For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ sub-tasks for each question digit before the “=” token. We created a framework for declaring n-digit algorithm hypotheses and testing them against the sub-tasks found in each model, mapping the results (see Table 2 for an example).
Our seven 2-layer addition models all implement our addition algorithm. Given their six nines accuracy and implementation of the same algorithm, we can confirm these models as known-good.
@@ 4.6 Investigating n-digit Mixed models
Table 3: Mixed models re-use most inserted addition-model nodes. Many inserted nodes become polysemantic during training - performing addition, positive-answer subtraction and negative-answer subtraction sub-tasks simultaneously. For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
@@ 4.8 Mixed model summary
The 7 mixed experimental models achieve five or six nines accuracy and contain the same sub-tasks, implying a common algorithm. However, without fully understanding this algorithm, we can’t confirm these models as known-good.
@@ 5 Conclusion
We successfully trained and verified known-good 5-, 6- and 10-digit, 2-layer, 3-head addition models implementing the same algorithm with minor variations.
We demonstrated component reuse by integrating an existing addition model into a larger “mixed" model for both addition and subtraction, achieving six nines accuracy. This integration helped us understand the mixed model’s algorithm. The mixed model reuses most inserted addition nodes, upgrading many to become polysemantic - performing both addition and subtraction subtasks simultaneously.
Our work supports (Michaud et al., 2023)’s assertion that many natural prediction problems can be broken into finite "quanta" computations essential for loss minimization. It also aligns with the idea that understanding a network’s functionality involves identifying and comprehending its sub-quanta.
@@ 6 Limitations and Future Work
Our research suggests several future directions. A key goal is to complete reverse-engineering the mixed model’s algorithm, understanding polysemantic nodes and OPR+SGN node output representations in the residual stream, and how MLP layers use this to predict answer digits accurately.
We aim to develop a comprehensive, known-good model for n-digit addition, subtraction, and multiplication. Our approach also could be applied to create known-good models in logical reasoning and planning.
Using smaller models to enhance LLM trustworthiness and capabilities is an active research area, including model composition (Bansal et al., 2024), LM up-scaling which essentially emulates fine-tuning a large pre-trained model using a small model (Mitchell et al., 2023), and inserting accurate models into untrained ones. Research on “spare” neurons in LLMs ((Voita et al., 2023; Hu et al., 2021)) suggests potential for small-scale modifications to fix erroneous circuits. Further exploration of using known-good models to improve LLMs is warranted.
Developing methods to incorporate compact known-good models into LLMs could democratize AI Safety research, allowing small teams to focus on specific areas and create quality components to improve LLMs.
Our automated framework for discovering algorithm sub-tasks in models accelerated our research and can test whether known-good components retain functionality when integrated into LLMs. While some aspects are specific to our math models, much is generic and reusable. Extending the framework to integrate output from complementary tools (e.g. ACDC’s node connectivity strength data (Conmy et al., 2023)) would provide more automated insights.
We progressed on a declarative language to describe algorithms in terms of necessary sub-tasks and a framework to test these descriptions against specific models. Further work is needed to fully realize its potential.
@@ Appendix E Appendix: Experimental models
Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Subtract Fails / M Heads used MLPs used
@@
5 1 3 30K 372001 9.4e-2 12621 N/A 15 6
5 2 3 15K 372001 1.6e-8 0 N/A 30 16
5 2 3 40K 372001 2.0e-9 0 N/A 22 15
6 2 3 15K 372001 1.7e-8 2 N/A 31 17
6 2 3 20K 173289 1.5e-8 0 N/A 28 17
6 2 3 20K 572091 7.0e-9 0 N/A 35 17
6 2 3 40K 372001 2.0e-9 0 N/A 29 17
10 2 3 40K 572091 7.0e-9 0 N/A 44 28
@@
6 2 3 40K 572091 2.4e-8 0 5 57 21
6 3 3 40K 572091 1.8e-8 0 3 70 35
6 3 3 80K 572091 1.6e-8 0 3 75 35
6 3 4 40K 372001 8.0e-9 0 0 72 26
6 3 4 40K 173289 1.4e-8 3 2 60 29
6 3 4 50K 572091 2.9e-8 0 4 79 29
10 3 3 50K 572091 6.3e-7 6 7 90 45
@@
6 4 3 40K 372001 3.0e-4 17 3120 115 53
Table 6: Main experimental models studied. The number of addition and subtraction failures per million questions is shown. The best 5-, 6- and 10-digit models are bolded.